a body of 0.02 kg falls from height 5m into a pile of sand.the body penetrtes the sand at ditance of 5m before stopping.what force has the sand exerted on the body

Grade:12

1 Answers

Arun
6 years ago
energy gained from gravity= mg*(10)
the sand has to provide the work..
force*distance = mg*10
forceaverage = mass*g*10/5 N
= 0.02 * 10 * 2 = 0.4 N